OpenStudy (anonymous):

How do I work out questions like this? State whether the three given points are collinear: a) (2,3) (4,4) (10,7) Thanks

OpenStudy (anonymous):

Find slope between first two points, slope between next two points. if they are same ; collinear

OpenStudy (anonymous):

4 - 3 / 4 - 2 = 1/2 = 0.5?

OpenStudy (anonymous):

then 7 - 4 / 10 - 4 = 3/6 = 1/2 = 0.5?

OpenStudy (anonymous):

Yeah, then colinear

OpenStudy (anonymous):

imranmeah91 do you want to take a online quiz for me?

OpenStudy (anonymous):

lol what????

OpenStudy (anonymous):

Only if you take my final

OpenStudy (anonymous):

Thank you for the explanation it helped !

OpenStudy (anonymous):

which final